If the price of a certain prescription medication increased from $25 to $50 and the number purchased decreased from 250 per week to 240 per week, what would the price elasticity of demand equal? (absolute value)

A. 0.06

B. 0.52

C. 1.92

D. 16.33
E. None of the above
Answer #1

Answer
Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=240-250=-10
average quantity=(240+250)/2=245
change in price=50-25=25
average price=(50+25)/2=37.5
Elasticity of demand=(-10/245)/(25/37.5)
=-0.0612244898
=-0.06
=0.06 (absolute value)

the elastcity od demand is 0.06

option A